A politician believes that 25% of all macroeconomists in senior positions would strongly support a proposal he wishes to advance. Suppose that this belief is correct and that five senior macroeconomists are approached at random. What is the probability that all 5 will strongly support the proposal?


1
Expert's answer
2022-02-14T16:24:15-0500

"n = 5\\

p = 0.25"

"P(X=x) = \binom{n}{x}p^x(1-p)^{n-x}"

We shall compute the probability that all five macroeconomists will support the proposal.

"p(X=5)= \binom{5}{5}\times 0.25^5 \times 0.75^{0}=0.25^5

= 0.001"

Therefore, the probability that all the five senior macroeconomists will support the proposal is 0.001